# Skew brace extensions, second cohomology and complements

### v1

@@ 4. Complements and supplements
Without loss of generality, assume that $\Phi(E)\neq E$. Suppose, for the sake of contradiction, that $I\not\subseteq\Phi(E)$. Then there exists an element $x\in I\setminus\Phi(E)$. Thus, we can extend $\{x\}$ to a generating set $S=\{g_{1},\ldots,g_{r-1},x\}$ of $E$, but no proper subset of $S$ generates $E$. It is sufficient prove that $\pi(S)=\{h_{1}:=\pi(g_{1}),\ldots,h_{r-1}:=\pi(g_{r-1}),\pi(x)=0\}$ generates $H$, as it will contradict the given supposition that $H$ is minimally generated by $r$ elements.
@@
For any two skew left braces $E_{1}$ and $E_{2}$, let $E:=E_{1}\oplus E_{2}$ be generated by a subset $X$ of $E$. Then it is not difficult to see that $E_{1}=\left<\pi_{1}(X)\right$> and $E_{2}=\left<\pi_{2}(X)\right$>, where $\pi_{i}:E\to E_{i}$, $i=1,2$, are the natural projections of skew braces. Now we prove
@@
which, considering the fact that $z=\big(\pi_{1}(z),\pi_{2}(z)\big)$ for all $z\in X$, is an image of the word $\epsilon w_{1}w_{2}$ on the elements of $X$. Thus $(x,y)\in E(X)$ and therefore $E=\left<X\right$>. The proof is complete. ∎
@@
Let $E$ be a cyclic skew brace of order not equal to any prime. Then $E$ admits a non-zero proper sub-skew brace ([4, Theorem A]), and therefore $E$ admits a unique maximal sub-skew brace equal to $\Phi(E)$. So if $E$ is a such a cyclic (cocyclic) skew brace, then $\Phi(E)\leq\Phi(E,+)$ ($\Phi(E)\leq\Phi(E,\circ)$). But, in general, there does not seem any clear connection. It will be interesting to explore relations among $\Phi(E)$, $\Phi(E,\circ)$ and $\Phi(E,+)$.
@@ 5. Reductions for split extension
In this section we study several reduction results for split extensions of finite skew braces by abelian groups (viewed as trivial brace). These results are brace analogs of group theory results from [16]. The following basic result is well known.
@@
Let $G$ be a group, and let $X$, $Y$ and $T$ be subgroups of $G$ such that $T$ is contained $X$. Then
@@
Let $I$ be a finite skew brace such that both $(I,+)$ and $(I,\circ)$ are nilpotent groups. Then it follows from [8, Corollary 4.3] that Sylow subgroups of $(I,+)$ are ideals of $I$. Furthermore, if $P_{i}$, $1\leq i\leq k$, are all distinct Sylow subgroups of $(I,+)$, then $I=P_{1}\oplus P_{2}\oplus\cdots\oplus P_{k}$. With this setting, the preceding theorem gives
@@
If $M$ is an $H$-submodule of $I$, then $M\times M$ is a $\Lambda_{H}$-submodule of $I\times I$.
@@
Let a finite abelian group $I$ be an $H$-module. Then there exists an $H$-composition series for $I$, and any two such composition series have the same length. Moreover, their composition factors are the same up to isomorphism and rearrangement.
@@
Let $I$ be a finite abelian group which is an $H$-module. Then any series of $H$-modules can be refined to an $H$-composition series.
@@
Let $E$ be a brace extension of a group $H$ by a finite abelian $p$-group II and
@@
Let $|I|=p^{n}$. Define $I^{p^{i}}:=\left<z^{p^{i}}\mid z\in I\right$>. It is well known that $I^{p^{i}}$ is a characteristic subgroup of $I$ for all $i\geq 1$ and $I^{p^{i}}/I^{p^{i+1}}$ is elementary abelian $p$-group. Thus $I^{p^{i}}$ is an ideal of $E$, and therefore an $H$-module. Since $I$ is finite, there exists $k\geq 1$ such that $I^{p^{r}}=\{0\}$. We get a series
@@
$x\in B_{1}\cap I=I_{1}\Rightarrow x\in B_{2}\cap I_{1}=I_{2}\Rightarrow\cdots\Rightarrow x\in B_{r}\cap I_{r-2}=I_{r-1}.$
@@
Acknowledgement: The first-named author is thankful to ISI Delhi for providing a good research ambiance. He thanks HRI Prayagraj for hosting him for two weeks, during which the work got impetus via personal interaction with the other author. He also thanks Arpan Karnar for the fruitful discussions. The second-named author acknowledges the partial support of DST-SERB Grant MTR/2021/000285.

### v2

@@ 4. Complements and supplements
Without loss of generality, assume that $\Phi(E)\neq E$. Suppose, for the sake of contradiction, that $I\not\subseteq\Phi(E)$. Then there exists an element $x\in I\setminus\Phi(E)$. Thus, we can extend $\{x\}$ to a generating set $S=\{g_{1},\ldots,g_{r-1},x\}$ of $E$, but no proper subset of $S$ generates $E$. It is sufficient to prove that $\pi(S)=\{h_{1}:=\pi(g_{1}),\ldots,h_{r-1}:=\pi(g_{r-1}),\pi(x)=0\}$ generates $H$, as it will contradict the given supposition that $H$ is minimally generated by $r$ elements.
@@
For any two skew braces $E_{1}$ and $E_{2}$, let $E:=E_{1}\oplus E_{2}$ be generated by a subset $X$ of $E$. Then it is not difficult to see that $E_{1}=\left<\pi_{1}(X)\right$> and $E_{2}=\left<\pi_{2}(X)\right$>, where $\pi_{i}:E\to E_{i}$, $i=1,2$, are the natural projections of skew braces. Now we prove
@@
Note that, in $E$, $(x,0)$ is a word in $x_{i_{1}},\ldots,x_{i_{k}}$ and $(0,y)$ is a word in $x_{j_{1}},\ldots,x_{j_{l}}$.
@@
which, considering the fact that $z=\big(\pi_{1}(z),\pi_{2}(z)\big)$ for all $z\in X$, is an image of the word $\epsilon w_{1}w_{2}$ on the elements of $X$, where $\epsilon=+$. Thus $(x,y)\in E(X)$ and therefore $E=\left<X\right$>. The proof is complete. ∎
@@
Let $E$ be a cyclic skew brace of order not equal to any prime. Then $E$ admits a non-zero proper sub-skew brace ([3, Theorem A]), and therefore $E$ admits a unique maximal sub-skew brace equal to $\Phi(E)$. So if $E$ is such a cyclic (cocyclic) skew brace, then $\Phi(E)\leq\Phi(E,+)$ ($\Phi(E)\leq\Phi(E,\circ)$). But, in general, there does not seem any clear connection. It will be interesting to explore relations among $\Phi(E)$, $\Phi(E,\circ)$ and $\Phi(E,+)$.
@@ 5. Reductions for split extension
In this section we study several reduction results on split extensions of finite skew braces by abelian groups (viewed as a trivial skew brace of abelian type). These results are brace analogs of group theory results from [13]. The following basic result is well known.
@@
Let $G$ be a group, and let $X$, $Y$ and $T$ be subgroups of $G$ such that $T$ is contained in $X$. Then
@@
Let $I$ be a finite skew brace such that both $(I,+)$ and $(I,\circ)$ are nilpotent groups. Then it follows from [7, Corollary 4.3] that Sylow subgroups of $(I,+)$ are ideals of $I$. Furthermore, if $P_{i}$, $1\leq i\leq k$, are all distinct Sylow subgroups of $(I,+)$, then $I=P_{1}\oplus P_{2}\oplus\cdots\oplus P_{k}$. With this setting, the preceding theorem gives
@@
Let $H$ be a skew brace. If $M$ is an $H$-submodule of an $H$-module II, then $M\times M$ is a $\Lambda_{H}$-submodule of $I\times I$.
@@
Let $H$ be a skew brace and $I$ be a finite $H$-module. Then there exists an $H$-composition series for $I$, and any two such composition series have the same length. Moreover, their composition factors are the same up to isomorphism and rearrangement.
@@
Let $H$ be a skew brace and $I$ be a finite $H$-module. Then any series of $H$-modules can be refined to an $H$-composition series.
@@
Let $E$ be a brace extension of a trivial skew brace $H$ by a finite abelian $p$-group II and
@@
Let $|I|=p^{n}$. Define $I^{p^{i}}:=\left<z^{p^{i}}\mid z\in I\right$>. It is well known that $I^{p^{i}}$ is a characteristic subgroup of $I$ for all $i\geq 1$ and $I^{p^{i}}/I^{p^{i+1}}$ is elementary abelian $p$-group. Thus $I^{p^{i}}$ is an ideal of $E$, and therefore an $H$-module. Since $I$ is finite, there exists $k\geq 1$ such that $I^{p^{k}}=\{0\}$. We get a series
@@
$x\in B_{1}\cap I=I_{1}\Rightarrow x\in B_{2}\cap I_{1}=I_{2}\Rightarrow\cdots\Rightarrow x\in B_{r-1}\cap I_{r-2}=I_{r-1}.$
@@
Acknowledgement: The first-named author is thankful to ISI Delhi for providing a good research ambience. He thanks HRI Prayagraj for hosting him for two weeks in September 2025, during which the work got impetus via personal interaction with the other author. He also thanks Arpan Karnar for fruitful discussions. The second-named author acknowledges the partial support of DST-SERB Grant MTR/2021/000285.